jayden490:

A bagel shop sells coffee in a container shaped like a rectangular prism. A graphic designer who works for the bagel shop drew the net below to create a design for the container. A Square in the middle where rectangle shapes are added around with the top length as 24cm and the side as 10 cm. The right side rectangle width is 24cm and the width is 17cm. The bottom side is marked as  10cm.

MakaylaChuck23:

Based on your description, the coffee container is shaped like a rectangular prism, and the net consists of a square base in the center with four rectangles attached around it. Here's what we can determine from the given dimensions: The top length is 24 cm, which means the width of the container is 24 cm. The side length is 10 cm, meaning the depth of the container is 10 cm. The right side rectangle has a width of 24 cm and a height of 17 cm, meaning the height of the container is 17 cm. Dimensions of the rectangular prism: Width: 24 cm Depth: 10 cm Height: 17 cm Finding the Surface Area: The surface area of a rectangular prism is given by: Surface Area = 2 𝑙 𝑤 + 2 𝑙 ℎ + 2 𝑤 ℎ Surface Area=2lw+2lh+2wh Plugging in the values: = 2 ( 24 × 10 ) + 2 ( 24 × 17 ) + 2 ( 10 × 17 ) =2(24×10)+2(24×17)+2(10×17) = 2 ( 240 ) + 2 ( 408 ) + 2 ( 170 ) =2(240)+2(408)+2(170) = 480 + 816 + 340 =480+816+340 = 1636  cm 2 =1636 cm 2 So, the total surface area of the coffee container is 1,636 cm². Would you like help with anything else related to this?